Solve the proportion 12 over 42 equals x over 14.. What is the value of x?

Answer by oscargut(2103)

Solve the proportion 12 over 42 equals x over 14.. What is the value of x?

12/42 = x/14
12*14 = 42x
42x = 168
x = 4
Answer: x = 4
You can ask me more at: mthman@gmail.com
Thanks